# Science Opportunities of Wet Extreme Mass-Ratio Inspirals

### v2

@@ II.3 Calibrating other Methods of Probing Black Hole Mass and Spin
The active type I fraction of galaxies with $5.5<\log_{10}(M_{\bullet})<7.0$ was estimated in a two-step process. First, following [99], we adopt the active ($\lambda_{\rm Edd}>10^{-3}$) fraction of $\sim 15-40$ % for galaxies with $5.5<\log_{10}(M_{\bullet})<6.5$, where $\lambda_{\rm Edd}$ is the Eddington ratio. This is broadly consistent with the active fraction of $\sim 24-34$ % for the same BH mass range estimated by [100], although a lower Eddington ratio is adopted. For a larger BH mass range $6.5<\log_{10}(M_{\bullet})<10.5$, the active ($\lambda_{\rm Edd}>10^{-3}$) fraction is $\sim 10-16$ % [101]. Furthermore, [102] yielded an active fraction of $\sim 10-22$ % for the host galaxy stellar mass range $9.0<\log_{10}(M_{\rm\star,~host})<10.0$. Assuming a $M_{\bullet}-M_{\rm\star,~host}$ relation (e.g., Eqn. 39 of [99]), this translates to a BH mass range of $5.0<\log_{10}(M_{\bullet})<6.0$.
Second, the active fraction of [99] should be multiplied by the type I fraction of active galaxies. The type I fraction of active galaxies ranges from $\sim 7$ % [103] to $\sim 20$ % [104, 105]. Therefore, the active type I fraction of galaxies with $5.5<\log_{10}(M_{\bullet})<6.5$ is $\sim 1-8$ % [99]. This is comparable to that of $\sim 4$ % for a larger BH mass range $6.5<\log_{10}(M_{\bullet})<10.5$ [101]. Hence, a few percent of galaxies with $5.5<\log_{10}(M_{\bullet})<7.0$ might be used to calibrate traditional X-ray measurements of the SMBH spin and optical measurements of the SMBH mass using precise GW spin and mass measurement.
@@
For X-ray counterparts of such wet EMRIs in type I AGNs, the most commonly used method to estimate BH spin is X-ray reflection spectroscopy [114, 115]. The reflection spectroscopy takes advantage of the skewed line profile of the 6.4 keV Fe line, which is featured with a sharp blueshifted peak due to the relativistic beaming effect and an extended redshifted wing due to the gravitational redshift of matter close to the innermost stable circular orbit (ISCO). More technical details, as well as descriptions of other techniques, can be found in the recent review on the observational constraints on BH spin [116].
@@
$L_{\rm 2-10~keV}=1.3\times 10^{44}{\rm~erg~s^{-1}}\left(\frac{M_{\bullet}}{10^{6}~M_{\odot}}\right)\left(\frac{\lambda_{\rm Edd}}{\kappa_{\rm 2-10~keV}}\right)$ (25)
@@
Table 3 presents a catalog of selected SMBHs with masses below $10^{7}M_{\odot}$, whose masses and spins have been determined using EM observations [123, 116, 124]. The table also includes SNRs for GW detections from LISA, TianQin, and Taiji, assuming a source with the same intrinsic parameters placed at the redshift of $z=0.3$ and with four years of observation. Although EM observations of SMBH spin typically carry significant uncertainties due to the limitations of current methods, GW observations from wet EMRIs provide far greater precision, with uncertainties as low as $10^{-6}\sim 10^{-4}$ for mass and $10^{-6}\sim 10^{-3}$ for spin [2, 68]. This comparison highlights the potential of wet EMRIs to calibrate traditional EM techniques, which rely on indirect methods such as X-ray reflection or virial mass estimators, often affected by large scatter and systematic biases. By integrating EM and GW data, Table 3 illustrates the critical role of wet EMRIs in improving the accuracy and reliability of SMBH mass and spin measurements, especially in type I AGNs.
@@ II.4 Testing Accretion Disk and Jet Models
Figure 2: Posterior distribution of the inclination angle $\iota$, shown for the dimensionless semi-latus rectum at $p_{0}\sim 8M_{\bullet}$ (left panel), and after tracing the evolution back to $p\sim 150M_{\bullet}$ due to GW radiation, and further to $p=410M_{\bullet}$ using Eqn. 14 with $\beta$-disk model (right panel). The orange vertical lines indicate the mean value along with the $1-\sigma$ uncertainty.
Figure 3: The relationship among the orbital angular momentum of the sBH $\hat{L}$, the SMBH spin direction $\hat{a}$, and the jet angle. The disk orientation $\hat{L}_{\rm disk}$ is inferred by tracing $\hat{L}$ backward from the LISA band ($p_{0}\sim 8M_{\bullet}$) to $p\sim 410M_{\bullet}$ (decoupling radius), where the sBH resides within the disk.
Figure 4: The uncertainty in the Lense-Thirring precession angle at $p\sim 150M_{\bullet}$, traced back from a semi-latus rectum of $p_{0}\sim 8M_{\bullet}$.
@@
To estimate the disk orientation at larger radii, we perform a backward evolution of the orbital trajectory—driven solely by GW damping and neglecting other dynamical interactions for simplicity—from the LISA band ($p\sim 8M_{\bullet}$) to $p\sim 150M_{\bullet}$. We then evolve the system further to $p=410M_{\bullet}$ using Eqn. 14 with the $\beta$-disk model. Under this assumption, the orbital angular momentum vector of the sBH, reconstructed at $p\sim 410M_{\bullet}$, serves as a proxy for the accretion disk orientation at that radius. This radius is significantly larger than the warp radius $R_{\rm warp}$ (see Eqn. 11) and corresponds to the outer region of the accretion disk, where disk forces are expected to dominate over relativistic effects, as discussed in Sec. II.2.
@@
The inclination angle $\iota$ of the sBH’s orbit, defined as [130]
@@ II.5.1 Wet EMRIs as bright sirens
Wet EMRIs, which involve compact objects spiraling into SMBHs within AGN disks, are ideal candidates for joint GW-EM observations. Similar to past events such as GW170817, wet EMRIs can act as bright sirens if their host AGNs are uniquely identified, allowing for precise redshift measurements. In addition, wet EMRIs have several advantages over other sources, such as binary neutron star or neutron star–black hole mergers. These advantages include higher expected detection rates, the ability to probe greater cosmological distances, and robust EM counterparts. Their relatively high localization precision and strong association with AGNs allow for a confident identification of host galaxies and their corresponding redshifts, making them powerful tools for constraining $H_{0}$ and other cosmological parameters at cosmological scales.
@@
Figure 5: The 1-$\sigma$ relative uncertainty of the Hubble parameter $H_{0}$ as a function of the number of bright EMRIs is shown for two representative wet EMRI models. As expected, the uncertainty on $H_{0}$ decreases with increasing numbers of bright EMRIs, reaching $1\%$ with seven bright sirens. Other simulated wet EMRI models exhibit similar behavior. The colored dots represent the calculated data points
@@
The inclusion of wet EMRI bright sirens in cosmological studies enhances the scope of GW-based standard siren measurements, complementing those from massive black hole binaries (MBHBs), binary neutron stars, and binary black holes. In parallel, MBHBs detected by space-based interferometers like LISA also serve as powerful bright sirens. As demonstrated by Tamanini et al. [146], LISA can constrain $H_{0}$ to sub-percent precision, particularly in its most optimistic configurations, with uncertainties reaching as low $0.5\%$. Importantly, MBHB standard sirens extend to redshifts as high as $z\sim 8$, significantly beyond the range accessible to wet EMRIs and EM probes. This broader redshift coverage allows MBHBs to provide independent and complementary constraints on cosmological parameters.
@@ II.5.2 Wet EMRIs as dark sirens
To statistically associate host galaxies with the GW source, the redshift prior $p(z_{\rm gw}\,|\,\mathbf{\Theta},\mathcal{H},I)$ is constructed by summing over all potential hosts - specifically, the host AGN galaxies in this study - within the GW localization volume:
@@ III Discussion and Conclusion
Second, in Sec. II we have made the implicit assumption that wet EMRIs are associated with visible AGNs. We have tried to include an estimation of the “dark” sources because of dust absorption, especially in the Hubble parameter measurement. There is, however, another scenario that the EMRI is brought to the vicinity of the massive BH because of disk migration, but the disk becomes quiescent during the time that this EMRI is observed by space-borne GW detectors. This scenario mainly affects wet EMRIs formed at the end of disk life cycles. If the lifetime of the disk is on average $T_{\rm disk}$ and the wet EMRI formation rate for that disk-assisted system is approximately once per $T_{\rm EMRI}$ (see the discussion in [22]), then this effect may affect a fraction of $T_{\rm EMRI}/T_{\rm disk}$ wet EMRIs. A more detailed study of this effect may include the time-dependent formation rate of wet EMRIs as the system starts to deplete sBHs within the nuclear star cluster.

### v3

@@ II.3 Calibrating other Methods of Probing Black Hole Mass and Spin
The active type I fraction of galaxies with $5.5<\log_{10}(M_{\bullet})<7.0$ was estimated in a two-step process. First, following Ref. [28], we adopt the active ($\lambda_{\rm Edd}>10^{-3}$) fraction of $\sim 15\%-40\%$ for galaxies with $5.5<\log_{10}(M_{\bullet})<6.5$, where $\lambda_{\rm Edd}$ is the Eddington ratio. This is broadly consistent with the active fraction of $\sim 24\%-34\%$ for the same BH mass range estimated by Ref. [49], although a lower Eddington ratio is adopted. For a larger BH mass range $6.5<\log_{10}(M_{\bullet})<10.5$, the active ($\lambda_{\rm Edd}>10^{-3}$) fraction is $\sim 10\%-16\%$ [5]. Furthermore, Ref. [101] yielded an active fraction of $\sim 10\%-22\%$ for the host galaxy stellar mass range $9.0<\log_{10}(M_{\rm\star,~host})<10.0$. Assuming a $M_{\bullet}-M_{\rm\star,~host}$ relation (e.g., Eq. 39 of Ref. [28]), this translates to a BH mass range of $5.0<\log_{10}(M_{\bullet})<6.0$.
Second, the active fraction of Ref. [28] should be multiplied by the type I fraction of active galaxies. The type I fraction of active galaxies ranges from $\sim 7\%$ [96] to $\sim 20\%$ [88, 100]. Therefore, the active type I fraction of galaxies with $5.5<\log_{10}(M_{\bullet})<6.5$ is $\sim 1\%-8\%$ [28]. This is comparable to that of $\sim 4\%$ for a larger BH mass range $6.5<\log_{10}(M_{\bullet})<10.5$ [5]. Hence, a few percent of galaxies with $5.5<\log_{10}(M_{\bullet})<7.0$ might be used to calibrate traditional X-ray measurements of the SMBH spin and optical measurements of the SMBH mass using precise GW spin and mass measurement.
@@
For X-ray counterparts of such wet EMRIs in type I AGNs, the most commonly used method to estimate BH spin is X-ray reflection spectroscopy [133, 50]. The reflection spectroscopy takes advantage of the skewed line profile of the 6.4 keV Fe line, which is featured with a sharp blueshifted peak due to the relativistic beaming effect and an extended redshifted wing due to the gravitational redshift of matter close to the innermost stable circular orbit. More technical details, as well as descriptions of other techniques, can be found in the recent review on the observational constraints on BH spin [112].
@@
$L_{\rm 2-10~keV}=1.3\times 10^{44}{\rm~erg~s^{-1}}\left(\frac{M_{\bullet}}{10^{6}~M_{\odot}}\right)\left(\frac{\lambda_{\rm Edd}}{\kappa_{\rm 2-10~keV}}\right)\ ,$ (25)
@@
Table 3 presents a catalog of selected SMBHs with masses below $10^{7}M_{\odot}$, whose masses and spins have been determined using EM observations [111, 112, 99]. The table also includes SNRs for GW detections from LISA, TianQin, and Taiji, assuming a source with the same intrinsic parameters placed at the redshift of $z=0.3$ and with four years of observation. Although EM observations of SMBH spin typically carry significant uncertainties due to the limitations of current methods, GW observations from wet EMRIs provide far greater precision, with uncertainties as low as $10^{-8}\sim 10^{-6}$ for mass and $10^{-7}\sim 10^{-5}$ for spin [12, 44], as also shown in FIG. 1. This comparison highlights the potential of wet EMRIs to calibrate traditional EM techniques, which rely on indirect methods such as X-ray reflection or virial mass estimators, often affected by large scatter and systematic biases. By integrating EM and GW data, Table 3 illustrates the critical role of wet EMRIs in improving the accuracy and reliability of SMBH mass and spin measurements, especially in type I AGNs.
@@ II.4 Testing Accretion Disk and Jet Models
Figure 3: Posterior distribution of the inclination angle $\iota$, shown for the dimensionless semi-latus rectum at $p_{0}\sim 8M_{\bullet}$ (left panel), and after tracing the evolution back to $p\sim 150M_{\bullet}$ due to GW radiation, and further to $p=410M_{\bullet}$ using Eq. 14 with $\beta$-disk model (right panel). The orange vertical lines indicate the mean value along with the $1-\sigma$ uncertainty.
Figure 4: Relationship among the orbital angular momentum of the sBH $\hat{L}$, the SMBH spin direction $\hat{a}$, and the jet angle. The disk orientation $\hat{L}_{\rm disk}$ is inferred by tracing $\hat{L}$ backward from the LISA band ($p_{0}\sim 8M_{\bullet}$) to $p\sim 410M_{\bullet}$ (decoupling radius), where the sBH resides within the disk.
Figure 5: Uncertainty in the Lense-Thirring precession angle at $p\sim 150M_{\bullet}$, traced back from a semi-latus rectum of $p_{0}\sim 8M_{\bullet}$.
@@
To estimate the disk orientation at larger radii, we perform a backward evolution of the orbital trajectory—driven solely by GW damping and neglecting other dynamical interactions for simplicity—from the LISA band ($p\sim 8M_{\bullet}$) to $p\sim 150M_{\bullet}$. We then evolve the system further to $p=410M_{\bullet}$ using Eq. 14 with the $\beta$-disk model. Under this assumption, the orbital angular momentum vector of the sBH, reconstructed at $p\sim 410M_{\bullet}$, serves as a proxy for the accretion disk orientation at that radius. This radius is significantly larger than the warp radius $R_{\rm warp}$ (see Eq. 11) and corresponds to the outer region of the accretion disk, where disk forces are expected to dominate over relativistic effects, as discussed in Sec. II.2.
@@
The inclination angle $\iota$ of the sBH’s orbit is defined as [118]
@@
The initial eccentricity of the system is set to $e=0.01$, consistent with expectations for disk-driven (wet) EMRIs. While idealized wet EMRI models often assume perfectly circular orbits ($e=0$), recent studies suggest that AGN disk interactions do not fully circularize the orbit. In particular, turbulence and multi-body resonances within the disk can induce residual eccentricities in the range $e\sim 10^{-4}$ to $10^{-2}$ in the LISA band, depending on specific disk properties such as the surface density and coherence time [127]. The choice of $e=0.01$ thus reflects a realistic and observationally motivated value for wet EMRIs, rather than a purely idealized scenario.
Furthermore, this residual eccentricity has implications for parameter estimation and waveform modeling. As shown in FIG. 1, LISA’s eccentricity measurement precision is expected to reach $\sim 10^{-5}$, enabling the detection of such small but nonzero values. In this work, we do not incorporate eccentricity when propagating posterior samples of the inclination angle $\iota$ to larger radii (e.g., $p\sim 150$ and $410\,M_{\bullet}$) for estimating uncertainties in disk geometry. The eccentricity contributes weakly to this propagation compared to inclination and spin, but its inclusion ensures a more realistic assessment of EMRI dynamics in AGN environments.
@@ II.5.1 Wet EMRIs as bright sirens
Wet EMRIs, which involve compact objects spiraling into SMBHs within AGN disks, are ideal candidates for joint GW-EM observations. Similar to past events such as GW170817, wet EMRIs can act as bright sirens if their host AGNs are uniquely identified, allowing for precise redshift measurements. In addition, wet EMRIs have several advantages over other sources, such as binary neutron star (BNS) or neutron star–black hole mergers. These advantages include higher expected detection rates, the ability to probe greater cosmological distances, and robust EM counterparts. Their relatively high localization precision and strong association with AGNs allow for a confident identification of host galaxies and their corresponding redshifts, making them powerful tools for constraining $H_{0}$ and other cosmological parameters at cosmological scales.
To measure $H_{0}$ using wet EMRI bright sirens, the luminosity distance $d_{\rm L}$ obtained from GW signals is combined with the redshift $z$ of the host AGN. Redshift uncertainties from spectroscopic measurements are typically in the range $z\sim 0.001-0.01$, consistent with forecasts from DESI [3] and LSST [33]. Since almost all host-identified EMRIs (bright sirens) are expected to lie at $z<0.3$, where spectroscopic redshifts are more accurate, the influence of redshift uncertainty is negligible for our analysis.
@@
Figure 6: 1-$\sigma$ relative uncertainty of the Hubble parameter $H_{0}$ as a function of the number of bright EMRIs is shown for two representative wet EMRI models. As expected, the uncertainty on $H_{0}$ decreases with increasing numbers of bright EMRIs, reaching $1\%$ with seven bright sirens. Other simulated wet EMRI models exhibit similar behavior. The colored dots represent the calculated data points
@@
The inclusion of wet EMRI bright sirens in cosmological studies enhances the scope of GW-based standard siren measurements, complementing those from massive black hole binaries (MBHBs), BHSs, and binary black holes (BBHs). In parallel, MBHBs detected by space-based interferometers like LISA also serve as powerful bright sirens. As demonstrated by Tamanini et al. [130], LISA can constrain $H_{0}$ to sub-percent precision, particularly in its most optimistic configurations, with uncertainties reaching as low $0.5\%$. Importantly, MBHB standard sirens extend to redshifts as high as $z\sim 8$, significantly beyond the range accessible to wet EMRIs and EM probes. This broader redshift coverage allows MBHBs to provide independent and complementary constraints on cosmological parameters.
@@ II.5.2 Wet EMRIs as dark sirens
To statistically associate host galaxies with the GW source, the redshift prior $p(z_{\rm gw}\,|\,\mathbf{\Theta},\mathcal{H},I)$ is constructed by summing over all potential hosts – specifically, the host AGN galaxies in this study – within the GW localization volume, as follows:
@@ III Discussion and Conclusion
Second, in Sec. II we have made the implicit assumption that wet EMRIs are associated with visible AGNs. We have tried to include an estimation of the “dark” sources because of dust absorption, especially in the Hubble parameter measurement. There is, however, another scenario that the EMRI is brought to the vicinity of the massive BH because of disk migration, but the disk becomes quiescent during the time that this EMRI is observed by space-borne GW detectors. This scenario mainly affects wet EMRIs formed at the end of disk life cycles. If the lifetime of the disk is on average $T_{\rm disk}$ and the wet EMRI formation rate for that disk-assisted system is approximately once per $T_{\rm EMRI}$ (see the discussion in Ref. [105]), then this effect may affect a fraction of $T_{\rm EMRI}/T_{\rm disk}$ wet EMRIs. A more detailed study of this effect may include the time-dependent formation rate of wet EMRIs as the system starts to deplete sBHs within the nuclear star cluster.
@@
Data Availability— The data that support the findings of this article are openly available at https://github.com/gwlyu/wet_EMRIs.git.
Appendix A Derivation of Sky Localization Volume
A.1 Fisher-matrix (quadratic) approximation
In the main text, the detector response is modeled using two orthogonal TDI channels $\alpha\in\{I,II\}$. For stationary Gaussian noise, the log-likelihood for parameters $\bm{\lambda}=\{\lambda_{i}\}$ is
$\ln\mathcal{L}(\bm{\lambda})=-\frac{1}{2}\sum_{\alpha\in\{I,II\}}\left\langle s_{\alpha}-h_{\alpha}(\bm{\lambda}),\,s_{\alpha}-h_{\alpha}(\bm{\lambda})\right\rangle\ ,$ (32)
where $s_{\alpha}$ is the (mock) data, $h_{\alpha}(\bm{\lambda})$ is the waveform template, and $\langle\cdot,\cdot\rangle$ is the noise-weighted inner product defined in Eq. (7).
Let $\hat{\bm{\lambda}}$ denote the maximum-likelihood (ML) point and define offsets $\Delta\lambda_{i}\equiv\lambda_{i}-\hat{\lambda}_{i}$. Expanding each channel waveform to first order about $\hat{\bm{\lambda}}$ gives
$h_{\alpha}(\bm{\lambda})\simeq h_{\alpha}(\hat{\bm{\lambda}})+\sum_{i}\left.\frac{\partial h_{\alpha}}{\partial\lambda_{i}}\right|_{\hat{\bm{\lambda}}}\Delta\lambda_{i}\ ,$ (33)
where the sum runs over model parameters.
Defining the residual at the ML point in each channel,
[…]
Using Eq. (33), the residual away from the ML point is
$s_{\alpha}-h_{\alpha}(\bm{\lambda})\simeq r_{\alpha}-\sum_{i}\left.\frac{\partial h_{\alpha}}{\partial\lambda_{i}}\right|_{\hat{\bm{\lambda}}}\Delta\lambda_{i}\ .$ (35)
Substituting into Eq. (32) and expanding yields
$\ln\mathcal{L}(\bm{\lambda})$ $\simeq-\frac{1}{2}\sum_{\alpha}\left[\langle r_{\alpha},r_{\alpha}\rangle-2\sum_{i}\left\langle r_{\alpha},\left.\frac{\partial h_{\alpha}}{\partial\lambda_{i}}\right|_{\hat{\bm{\lambda}}}\right\rangle\Delta\lambda_{i}+\sum_{i,j}\left\langle\left.\frac{\partial h_{\alpha}}{\partial\lambda_{i}}\right|_{\hat{\bm{\lambda}}},\left.\frac{\partial h_{\alpha}}{\partial\lambda_{j}}\right|_{\hat{\bm{\lambda}}}\right\rangle\Delta\lambda_{i}\Delta\lambda_{j}\right]\ ,$ (36)
The linear term vanishes because the score is zero at the ML point, as follow:
$0=\left.\frac{\partial\ln\mathcal{L}}{\partial\lambda_{i}}\right|_{\hat{\bm{\lambda}}}=\sum_{\alpha}\left\langle r_{\alpha},\left.\frac{\partial h_{\alpha}}{\partial\lambda_{i}}\right|_{\hat{\bm{\lambda}}}\right\rangle\$. (37)
Subtracting the ML value $\ln\mathcal{L}(\hat{\bm{\lambda}})=-(1/2)\sum_{\alpha}\langle r_{\alpha},r_{\alpha}\rangle$ then gives the standard quadratic (Fisher) approximation,
$2\ln\frac{\mathcal{L}(\bm{\lambda})}{\mathcal{L}(\hat{\bm{\lambda}})}\simeq\sum_{i,j}\Delta\lambda_{i}\,\Gamma_{ij}\,\Delta\lambda_{j}\ ,$ (38)
where the Fisher information matrix is
$\Gamma_{ij}\equiv\sum_{\alpha\in\{I,II\}}\left\langle\left.\frac{\partial h_{\alpha}}{\partial\lambda_{i}}\right|_{\hat{\bm{\lambda}}},\left.\frac{\partial h_{\alpha}}{\partial\lambda_{j}}\right|_{\hat{\bm{\lambda}}}\right\rangle\$. (39)
If the prior is approximately flat across the region of significant likelihood, the posterior is locally Gaussian with covariance
$\Sigma_{ij}\simeq(\Gamma^{-1})_{ij}\ .$ (40)
The off-diagonal elements $\Sigma_{ij}$ encode covariances (and hence correlations) between parameters $\lambda_{i}$ and $\lambda_{j}$, $\mathrm{Cov}(\lambda_{i},\lambda_{j})=\Sigma_{ij}$, while the diagonal elements give the marginalized variances, $\mathrm{Var}(\lambda_{i})=\Sigma_{ii}\equiv\sigma_{\lambda_{i}}^{2}$.
A.2 Joint $p$-probability ellipsoid in $N$ dimensions
Consider an $N$-dimensional subset of parameters $\bm{\vartheta}$ (e.g. $\bm{\vartheta}=(\ln r,\theta,\phi)$) with covariance sub-matrix $\Sigma_{\bm{\vartheta}}$ (the corresponding sub-block of $\Sigma_{ij}$). Define the Mahalanobis quadratic form
$Q\equiv\Delta\bm{\vartheta}^{\mathsf{T}}\,\Sigma_{\bm{\vartheta}}^{-1}\,\Delta\bm{\vartheta}\ .$ (41)
If $\Delta\bm{\vartheta}$ is multivariate Gaussian with covariance $\Sigma_{\bm{\vartheta}}$, then
$Q\sim\chi^{2}_{N}\ ,$ (42)
The joint $p$-probability region is the ellipsoid
$\mathcal{E}_{p}\equiv\left\{\Delta\bm{\vartheta}:\;Q\leq A_{p,N}^{2}\right\}\ ,$ (43)
where $A_{p,N}^{2}$ is fixed by the $\chi^{2}_{N}$ quantile, as follow:
$p=P(\chi^{2}_{N}\leq A_{p,N}^{2})=\frac{\gamma\!\left(\frac{N}{2},\,\frac{A_{p,N}^{2}}{2}\right)}{\Gamma\!\left(\frac{N}{2}\right)}\ .$ (44)
Here $\gamma(\cdot,\cdot)$ and $\Gamma(\cdot)$ are the lower incomplete and complete $\Gamma$ functions (not to be confused with the Fisher matrix $\Gamma_{ij}$).
Equivalently, one may write
$A_{p,N}^{2}=\chi^{2}_{N,p}\ ,$ (45)
where $\chi^{2}_{N,p}$ denotes the $p$-quantile of the $\chi^{2}$ distribution with $N$ degrees of freedom. For the commonly used $p=0.6827$ joint credible regions,
$N=2:\quad$ $A_{0.68,\,2}^{2}=\chi^{2}_{2,\,0.68}\simeq 2.296\ ,\;A_{0.68,\,2}\simeq 1.515\ ,$ (46)
$N=3:\quad$ $A_{0.68,\,3}^{2}=\chi^{2}_{3,\,0.68}\simeq 3.527\ ,\;A_{0.68,\,3}\simeq 1.878\ .$ (47)
A.3 Hypervolume of the ellipsoid and numerical prefactors
To compute the hypervolume enclosed by Eq. (43), we apply a whitening transformation. Let $\Sigma_{\bm{\vartheta}}=LL^{\mathsf{T}}$ be a Cholesky factorization and define
$\mathbf{y}\equiv L^{-1}\Delta\bm{\vartheta}\ .$ (48)
Then,
$Q=\Delta\bm{\vartheta}^{\mathsf{T}}\Sigma_{\bm{\vartheta}}^{-1}\Delta\bm{\vartheta}=\mathbf{y}^{\mathsf{T}}\mathbf{y}=\|\mathbf{y}\|^{2}\ ,$ (49)
so $\mathcal{E}_{p}$ maps to the $N$-ball $\|\mathbf{y}\|\leq A_{p,N}$. The Jacobian determinant is
$\left|\frac{\partial\Delta\bm{\vartheta}}{\partial\mathbf{y}}\right|=|\det L|=\sqrt{\det\Sigma_{\bm{\vartheta}}}\ .$ (50)
Hence the $p$-probability ellipsoid hypervolume is
$V_{p}^{(N)}=V_{\mathrm{ball},N}(A_{p,N})\;\sqrt{\det\Sigma_{\bm{\vartheta}}}\ ,$ (51)
where the $N$-ball volume of radius $A$ is
$V_{\mathrm{ball},N}(A)=\frac{\pi^{N/2}}{\Gamma\!\left(\frac{N}{2}+1\right)}\,A^{N}\ .$ (52)
In particular, the $N$-ball volumes are
$V_{\mathrm{ball},2}(A)=\pi A^{2}\quad\text{(area)}\ ,\qquad V_{\mathrm{ball},3}(A)=\frac{4\pi}{3}A^{3}\quad\text{(volume)}\ .$ (53)
For the joint $p=0.6827$ region, this yields
$N=2:\;$ $V_{0.68}^{(2)}=\pi A_{0.68,2}^{2}\,\sqrt{\det\Sigma_{\bm{\vartheta}}}\;\simeq\;7.21\,\sqrt{\det\Sigma_{\bm{\vartheta}}}\ ,$ (54)
$N=3:\;$ $V_{0.68}^{(3)}=\frac{4\pi}{3}A_{0.68,3}^{3}\,\sqrt{\det\Sigma_{\bm{\vartheta}}}\;\simeq\;27.74\,\sqrt{\det\Sigma_{\bm{\vartheta}}}\ .$ (55)
A.4 Sky solid angle and 3D localization volume
We specialize the joint $p$-ellipsoid construction to (i) sky angles and (ii) the 3D localization volume.
Solid-angle (sky) uncertainty. Let $\bm{\vartheta}=(\theta,\phi)$ with covariance sub-matrix $\Sigma_{\theta,\phi}$ (the $2\times 2$ sub-block of $\Sigma_{ij}$ evaluated at $\hat{\bm{\lambda}}$). The area of the joint $p$ ellipse in the coordinate plane $(\theta,\phi)$ is
$V_{p}^{(2)}\;=\;\pi A_{p,2}^{2}\,\sqrt{\det\Sigma_{\theta,\phi}}\ ,\qquad A_{p,2}^{2}=\chi^{2}_{2,p}\ .$ (56)
The physical solid-angle element on the unit sphere is
$d\Omega=\sin\theta\,d\theta\,d\phi\ .$ (57)
If the uncertainty region is sufficiently small that $\sin\theta$ may be treated as constant across it, evaluated at the best-fit sky location $\hat{\theta}$, then the joint $p$ solid-angle uncertainty is
$\Delta\Omega_{p}\;\simeq\;\sin\hat{\theta}\;V_{p}^{(2)}\;=\;\pi A_{p,2}^{2}\,\sin\hat{\theta}\,\sqrt{\det\Sigma_{\theta,\phi}}\ .$ (58)
For $p=0.6827$, $A_{0.68,\,2}\simeq 1.515$
$\boxed{\Delta\Omega_{0.68}\simeq 7.21\,\sin\hat{\theta}\,\sqrt{\det\Sigma_{\theta,\phi}}}\ .$ (59)
Physical sky-localization volume. For 3D localization take $\bm{\vartheta}=(\ln r,\theta,\phi)$ with covariance sub-matrix $\Sigma_{\ln r,\theta,\phi}$. The physical volume element is
$dV=r^{2}\sin\theta\,dr\,d\theta\,d\phi=r^{3}\sin\theta\,d(\ln r)\,d\theta\,d\phi\ .$ (60)
[…]
$V_{p}^{(3)}\;=\;\frac{4\pi}{3}A_{p,3}^{3}\,\sqrt{\det\!\left(\Sigma_{\ln r,\theta,\phi}\right)}\ .\qquad A_{p,3}^{2}=\chi^{2}_{3,p}\ .$ (61)
Multiplying by the physical Jacobian factor from Eq. (60) gives the physical localization volume
$\Delta V_{p}\;\simeq\;\hat{r}^{\,3}\sin\hat{\theta}\;\frac{4\pi}{3}A_{p,3}^{3}\,\sqrt{\det\!\left(\Sigma_{\ln r,\theta,\phi}\right)}\ .$ (62)
For $p=0.6827$, $A_{0.68,\,3}\simeq 1.878$, hence
$\boxed{\Delta V_{0.68}\simeq 27.74\,\hat{r}^{\,3}\sin\hat{\theta}\,\sqrt{\det\!\left(\Sigma_{\ln r,\theta,\phi}\right)}}\ .$ (63)
This expression accounts for correlations among $(\ln r,\theta,\phi)$ through the determinant of the full $3\times 3$ covariance sub-matrix.
A.5 Remark on factorized approximations
A common shortcut is to “factorize” the 3D localization volume into a radial part and a sky part, as follows:
$\Delta V\approx\hat{r}^{3}\,\sigma_{\ln r}\,\Delta\Omega\ .$ (64)
This approximation is valid only if: (i) $\ln r$ is (nearly) uncorrelated with the sky angles $(\theta,\phi)$, so the 3D covariance is approximately block-diagonal; and (ii) $\sigma_{\ln r}$ and $\Delta\Omega$ are defined so that they correspond to the same joint probability content $p$.
In general the posterior in $(\ln r,\theta,\phi)$ is a correlated 3D Gaussian (under the Fisher approximation), so the correct joint $p$ region is a 3D ellipsoid. Its physical volume for $p=0.6827$ is expressed as Eq. (62). Therefore an expression such as
$\Delta V_{0.68}=\hat{r}^{3}\,\sigma_{\ln r}\,\pi\sin\hat{\theta}\,\sqrt{\det(\Sigma_{\theta,\phi})}\ .$ (65)
is generally inconsistent: it (a) ignores correlations between distance and sky position, and (b) mixes a 1D and 2D uncertainty in a way that does not represent a joint 68% region; the correct joint-3D numerical factor is $\simeq 28$, not $\pi$.